# Functional differential equations driven by càdlàg rough paths

### v2

@@ 3.1. Classical RDEs
where $y_{0}\in\mathbb{R}^{k}$, $f\in C^{3}_{b}(\mathbb{R}^{k};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ and $\mathbf{X}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ for $p\in(2,3)$. While the existence and uniqueness of solutions to the RDE (3.1) driven by a continuous rough path and the continuity of the solution map were first proven by Lyons [Lyo98], the analogous results for RDEs driven by càdlàg rough paths were more recently obtained by Friz and Zhang [FZ18]. As an application of Theorem 2.4 and Theorem 2.5, one can recover these results, demonstrating that Assumption 2.1 is, indeed, a natural generalization of the classical assumptions of the coefficients of a rough differential equation. Furthermore, note that Corollary 3.1 presents the continuity of the solution map with respect to the controlled path norm, which slightly generalizes [FZ18, Theorem 3.8].
@@
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the RDE (3.1). Moreover, let $\tilde{y}_{0}\in\mathbb{R}^{k}$, $\tilde{f}\in C^{3}_{b}(\mathbb{R}^{k};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$, $\tilde{\mathbf{X}}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ with corresponding solution $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}([0,T];\mathbb{R}^{d})$, and suppose that $\|Y,Y^{\prime}\|_{X,p},\|\tilde{Y},\tilde{Y}^{\prime}\|_{\tilde{X},p}\leq K$, for some ${K>0}$. Then, we have the estimate
    $|Y_{0}-\tilde{Y}_{0}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p}\lesssim|y_{0}-\tilde{y}_{0}|+\|f-\tilde{f}\|_{C^{2}_{b}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p},$
    where the implicit multiplicative constant depends on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$.
In order to apply the existence and uniqueness result presented in Theorem 2.4, and also the continuity result presented in Theorem 2.5, we need to check that the vector field $f$ in the RDE (3.1) satisfies Assumption 2.1. This is the content of the next lemma, which we formulate slightly more general, with regard to the dimensions of the underlying spaces, for later use.
@@
Let $f\in C^{3}_{b}(\mathbb{R}^{m};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ and $X,\tilde{X}\in D^{p}([0,T];\mathbb{R}^{d})$. The non-anticipative functional
@@
satisfies Assumption 2.1.
@@
Fix $(s,t)\in\Delta_{T}$ and let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$, $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}$.
@@
Lipschitz conditions. Fix $K>0$ and assume that $\|Y,Y^{\prime}\|_{X,p,[s,t]},\|\tilde{Y},\tilde{Y}^{\prime}\|_{\tilde{X},p,[s,t]}\leq K$. The proofs work verbatim as the proofs of [FZ18, Lemma 3.1 and Lemma 3.7]. The constant $C_{F,K,X,\tilde{X}}$ depends on $p$ and $\|f\|_{C^{3}_{b}}$, $K$, $\|X\|_{p}$ and $\|\tilde{X}\|_{p}$. ∎
@@
(ii) To apply the continuity result presented in Theorem 2.5, we need to ensure that the functionals satisfy the required assumptions. For $(F,F^{\prime}),(\tilde{F},\tilde{F}^{\prime})$, this is given in Lemma 3.2, and further,
$(F-\tilde{F},(F-\tilde{F})^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})),$
$((F-\tilde{F})(Y),(F-\tilde{F})^{\prime}(Y,Y^{\prime})):=(f(Y)-\tilde{f}(Y),\mathrm{D}f(Y)Y^{\prime}-\mathrm{D}\tilde{f}(Y)Y^{\prime}),$
satisfies Assumption 2.1 (i), since $C^{3}_{b}$ is a vector space. Thus we have, $C_{F-\tilde{F}}\lesssim_{p}\|f-\tilde{f}\|_{C^{2}_{b}}$. ∎
@@ 3.2. Controlled RDEs
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the controlled rough differential equation (3.2). Moreover, let $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, $(\tilde{\alpha},\tilde{\alpha}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{e})$, $\tilde{f}\in C^{3}_{b}(\mathbb{R}^{k+e};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$, with corresponding solution $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, and suppose that $\|Y,Y^{\prime}\|_{X,p},\|\tilde{Y},\tilde{Y}^{\prime}\|_{X,p}\leq K$, for some $K>0$. Then, we have the estimate
    $|Y_{0}-\tilde{Y}_{0}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,p}$
@@
In order to apply the existence and uniqueness result presented in Theorem 2.4, and also the continuity result in Theorem 2.5, we need to check that the vector field in the RDE (3.2) satisfies Assumption 2.1. This is the content of the next lemma. Note that it will be sufficient to check Assumption 2.1 (ii) for $X=\tilde{X}$, since we do not establish stability results with respect to the driving rough path in this subsection. (This also applies to the following subsections.)
@@
satisfies Assumption 2.1 for $X=\tilde{X}$.
@@
and it holds that $\|Y\|_{p,[s,t]}\leq\|Y\|_{p,[s,t)}+|Y_{t-,t}|$ as well as $\|Y\|_{p,[s,t]}\leq C_{p}(1+(|Y^{\prime}_{s}|+\|Y^{\prime}\|_{p,[s,t]})\cdot\|X\|_{p,[s,t]}+\|R^{Y}\|_{\frac{p}{2},[s,t]})$. Applying Lemma 3.2 to the enlarged controlled path $((\alpha,Y),(\alpha^{\prime},Y^{\prime}))$,
@@
$|(\alpha,Y)_{s}-(\alpha,\tilde{Y})_{s}|+\|(\alpha,Y)-(\alpha,\tilde{Y})\|_{p,[s,t]}=|Y_{s}-\tilde{Y}_{s}|+\|Y-\tilde{Y}\|_{p,[s,t]}$
@@
$\|(\alpha,Y),(\alpha^{\prime},Y^{\prime});(\alpha,\tilde{Y}),(\alpha^{\prime},\tilde{Y}^{\prime})\|_{X,p}=\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,p},$
@@
(ii) To apply the continuity result presented in Theorem 2.5, we need to ensure that the functionals satisfy the required assumptions. For $(F,F^{\prime}),(\tilde{F},\tilde{F}^{\prime})$, this is given in Lemma 3.4. Analogously, since $f-\tilde{f}\in C^{3}_{b}$, we note that for
@@
$(F-\tilde{F},F^{\prime}-\tilde{F}^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})),$
$((F-\tilde{F})(Y),(F-\tilde{F})^{\prime}(Y,Y^{\prime}))$
@@
satisfies Assumption 2.1 (i). Here, $C_{F-\tilde{F}}=\|f-\tilde{f}\|_{C^{2}_{b}}+|\alpha_{0}-\tilde{\alpha}_{0}|+\|\alpha-\tilde{\alpha}\|_{p}+\|\alpha,\alpha^{\prime};\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$ up to a multiplicative constant, which depends on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $\|\alpha\|_{p}$, $\|\tilde{\alpha}\|_{p}$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, $\|\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$. ∎
@@ 3.3. RDEs with discrete time dependence
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the RDE with time discrete dependence (3.3). Moreover, let $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ with corresponding solution $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, and suppose that $\|Y,Y^{\prime}\|_{X,p},\|\tilde{Y},\tilde{Y}^{\prime}\|_{X,p}\leq K$, for some ${K>0}$. Then, we have the estimate
    $|Y_{0}-\tilde{Y}_{0}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,p}\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p},$
@@
with $(\bar{Y},\bar{Y}^{\prime})=((Y,Y,\ldots,Y),(Y^{\prime},Y^{\prime},\ldots,Y^{\prime}))\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k(\ell+1)})$. It follows analogously to Lemma 3.2 that the functional satisfies Assumption 2.1 with constants depending additionally on $\ell$, that is, $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends on $p$ and $\ell$, and $C_{F,K,X,X}$ depends on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, and $\|X\|_{p}$. Note that it is sufficient to check Assumption 2.1 (ii) for $X=\tilde{X}$. We can thus apply Theorem 2.4 to show that there exists a unique solution to the RDE (3.3) on the interval $[0,r_{1})$. We now aim to solve the RDE (3.3) iteratively on the subintervals $[r_{i},r_{i+1})$, $i=1,\ldots,\ell$, with $r_{\ell+1}=T$. Given the solution on $[r_{i-1},r_{i})$, with $r_{0}=0$, the value $Y_{r_{i}}$ is determined by the jump of $\mathbf{X}$ at time $r_{i}$. We therefore consider $(y_{i},y_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}];\mathbb{R}^{k})$, where
@@
for $i=\ell$. Analogously to Lemma 3.4, we can show that the functional satisfies Assumption 2.1 with constants depending additionally on $\ell$, that is $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends on $p$ and $\ell$, see the definition of $(\alpha_{i},\alpha_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}])$, and $C_{F,K,X,X}$ depends on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, and $\|X\|_{p}$. Note that it is again sufficient to check Assumption 2.1 (ii) for $X=\tilde{X}$. We can thus again apply Theorem 2.4 to show that there exists a unique solution to the RDE (3.3) on the interval $[r_{i},r_{i+1})$, that is
@@
For $(F,F^{\prime})$, this is shown in the proof of (i), and as we aim to obtain continuity of the solution map as a function of the initial condition $(y,y^{\prime})$, not the vector field $f$, on the interval $[0,r_{1}]$ we may consider $(F,F^{\prime})=(\tilde{F},\tilde{F}^{\prime})$, so, $(F-\tilde{F},F^{\prime}-\tilde{F}^{\prime})=0$. Theorem 2.5 now gives that
$\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,r_{1}]}+\|R^{Y}-R^{\tilde{Y}}\|_{p,[0,r_{1}]}\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p,[0,r_{1}]},$
@@
Local estimate on $[r_{i},r_{i+1}]$, $i=1,\ldots,\ell$. To apply the continuity result presented in Theorem 2.5, we need to ensure that the functionals satisfy the required assumptions. For $(F,F^{\prime}),(\tilde{F},\tilde{F}^{\prime})$, this is shown in the proof of (i), and for $(F-\tilde{F},F^{\prime}-\tilde{F}^{\prime})$, in the proof of Corollary 3.3 (ii), where the constant $C_{F-\tilde{F}}$ depends additionally on $\ell$. Theorem 2.5 then implies that
$\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[r_{i},r_{i+1}]}+\|R^{Y}-R^{\tilde{Y}}\|_{p,[r_{i},r_{i+1}]}\lesssim|Y_{r_{i}}-\tilde{Y}_{r_{i}}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p,[r_{i},r_{i+1}]},$
@@
$\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2}}\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p},$

### v3

@@ 3.1. Classical RDEs
where $y_{0}\in\mathbb{R}^{k}$, $f\in C^{3}_{b}(\mathbb{R}^{k};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ and $\mathbf{X}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ for $p\in(2,3)$. While the existence and uniqueness of solutions to the RDE (3.1) driven by a continuous rough path and the continuity of the solution map were first proven by Lyons [Lyo98], the analogous results for RDEs driven by càdlàg rough paths were more recently obtained by Friz and Zhang [FZ18]. As an application of Theorem 2.3 and Theorem 2.5, one can recover these results, demonstrating that Assumption 2.1 and 2.4 are, indeed, natural generalizations of the classical assumptions of the coefficients of a rough differential equation. Furthermore, note that Corollary 3.1 presents the continuity of the solution map with respect to the controlled path norm, which slightly generalizes [FZ18, Theorem 3.8].
@@
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the RDE (3.1). Moreover, let $\tilde{y}_{0}\in\mathbb{R}^{k}$, $\tilde{f}\in C^{3}_{b}(\mathbb{R}^{k};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$, $\widetilde{\mathbf{X}}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ with corresponding solution $(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{\widetilde{X}}([0,T];\mathbb{R}^{d})$, and suppose that $\|Y,Y^{\prime}\|_{X,p},\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{\widetilde{X},p}\leq K$, for some ${K>0}$. Then, we have the estimate
    $|Y_{0}-\widetilde{Y}_{0}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,\widetilde{X},p}\lesssim|y_{0}-\tilde{y}_{0}|+\|f-\tilde{f}\|_{C^{2}_{b}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p},$
    where the implicit multiplicative constant depends on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\widetilde{\mathbf{X}}\|_{p}$.
In order to apply the existence and uniqueness result presented in Theorem 2.3, and also the continuity result presented in Theorem 2.5, we need to check that the vector field $f$ in the RDE (3.1) satisfies Assumption 2.1 and 2.4. This is the content of the next lemma, which we formulate slightly more general, with regard to the dimensions of the underlying spaces, for later use.
@@
Let $f\in C^{3}_{b}(\mathbb{R}^{m};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ and $X,\widetilde{X}\in D^{p}([0,T];\mathbb{R}^{d})$. The non-anticipative functional
@@
satisfies Assumption 2.4 (i) and (ii), and, in particular, Assumption 2.1, given $X,\widetilde{X}$.
@@
Fix $(s,t)\in\Delta_{T}$ and let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$, $(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{\widetilde{X}}$.
@@
Lipschitz conditions. Fix $K>0$ and assume that $\|Y,Y^{\prime}\|_{X,p,[s,t]},\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{\widetilde{X},p,[s,t]}\leq K$. The proofs work verbatim as the proofs of [FZ18, Lemma 3.1 and Lemma 3.7]. The constant $C_{F,K,X,\widetilde{X}}$ depends on $p$ and $\|f\|_{C^{3}_{b}}$, $K$, $\|X\|_{p}$ and $\|\widetilde{X}\|_{p}$. ∎
@@
(ii) To apply the continuity result presented in Theorem 2.5, we need to ensure that the functionals satisfy the required assumptions. For $(F,F^{\prime}),(\widetilde{F},\widetilde{F}^{\prime})$, this is given in Lemma 3.2, and further,
$(F-\widetilde{F},(F-\widetilde{F})^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})),$
$((F-\widetilde{F})(Y),(F-\widetilde{F})^{\prime}(Y,Y^{\prime})):=(f(Y)-\tilde{f}(Y),\mathrm{D}f(Y)Y^{\prime}-\mathrm{D}\tilde{f}(Y)Y^{\prime}),$
satisfies the corresponding estimates in Assumption 2.4, since $C^{3}_{b}$ is a vector space. Thus we have, $C_{F-\widetilde{F}}\lesssim_{p}\|f-\tilde{f}\|_{C^{2}_{b}}$. ∎
@@ 3.2. Controlled RDEs
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the controlled rough differential equation (3.2). Moreover, let $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, $(\tilde{\alpha},\tilde{\alpha}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{e})$, $\tilde{f}\in C^{3}_{b}(\mathbb{R}^{k+e};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$, with corresponding solution $(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, and suppose that $\|Y,Y^{\prime}\|_{X,p},\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p}\leq K$, for some $K>0$. Then, we have the estimate
    $|Y_{0}-\widetilde{Y}_{0}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p}$
@@
In order to apply the existence and uniqueness result presented in Theorem 2.3, and also the continuity result in Theorem 2.5, we need to check that the vector field in the RDE (3.2) satisfies Assumption 2.4. This is the content of the next lemma. Note that it will be sufficient to check Assumption 2.4 (ii) for $X=\widetilde{X}$, since we do not establish stability results with respect to the driving rough path in this subsection. (This also applies to the following subsections.)
@@
satisfies Assumption 2.4 given $X=\widetilde{X}$.
@@
and it holds that $\|Y\|_{p,[s,t]}\leq\|Y\|_{p,[s,t)}+|Y_{t-,t}|$ as well as $\|Y\|_{p,[s,t]}\leq C_{p}(1+(|Y^{\prime}_{s}|+\|Y^{\prime}\|_{p,[s,t]})\cdot\|X\|_{p,[s,t]}+\|R^{Y}\|_{\frac{p}{2},[s,t]})$. Applying Lemma 3.2 to the enlarged controlled path $((\alpha,Y),(\alpha^{\prime},Y^{\prime}))$, it follows that
@@
$|(\alpha,Y)_{s}-(\alpha,\widetilde{Y})_{s}|+\|(\alpha,Y)-(\alpha,\widetilde{Y})\|_{p,[s,t]}=|Y_{s}-\widetilde{Y}_{s}|+\|Y-\widetilde{Y}\|_{p,[s,t]}$
@@
$\|(\alpha,Y),(\alpha^{\prime},Y^{\prime});(\alpha,\widetilde{Y}),(\alpha^{\prime},\widetilde{Y}^{\prime})\|_{X,p}=\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p},$
@@
(ii) To apply the continuity result presented in Theorem 2.5, we need to ensure that the functionals satisfy the required assumptions. For $(F,F^{\prime}),(\widetilde{F},\widetilde{F}^{\prime})$, this is given in Lemma 3.4. Analogously, since $f-\tilde{f}\in C^{3}_{b}$, we note that for
@@
$(F-\widetilde{F},F^{\prime}-\widetilde{F}^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})),$
$((F-\widetilde{F})(Y),(F-\widetilde{F})^{\prime}(Y,Y^{\prime}))$
@@
satisfies the corresponding estimates in Assumption 2.4 with
$C_{F-\widetilde{F}}=\|f-\tilde{f}\|_{C^{2}_{b}}+|\alpha_{0}-\tilde{\alpha}_{0}|+\|\alpha-\tilde{\alpha}\|_{p}+\|\alpha,\alpha^{\prime};\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$
up to a multiplicative constant, which depends on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $\|\alpha\|_{p}$, $\|\tilde{\alpha}\|_{p}$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, $\|\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$. ∎
@@ 3.3. RDEs with discrete time dependence
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the RDE with time discrete dependence (3.3). Moreover, let $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ with corresponding solution $(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, and suppose that $\|Y,Y^{\prime}\|_{X,p},\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p}\leq K$, for some ${K>0}$. Then, we have the estimate
    $|Y_{0}-\widetilde{Y}_{0}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p}\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p},$
@@
with $(\bar{Y},\bar{Y}^{\prime})=((Y,Y,\ldots,Y),(Y^{\prime},Y^{\prime},\ldots,Y^{\prime}))\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k(\ell+1)})$. It follows analogously to Lemma 3.2 that the functional satisfies Assumption 2.4 (i) and (ii) with constants depending additionally on $\ell$, that is, $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends on $p$ and $\ell$, and $C_{F,K,X,X}$ depends on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, and $\|X\|_{p}$. Note that it is sufficient to check Assumption 2.4 (ii) for $X=\widetilde{X}$. We can thus apply Theorem 2.3 to show that there exists a unique solution to the RDE (3.3) on the interval $[0,r_{1})$. We now aim to solve the RDE (3.3) iteratively on the subintervals $[r_{i},r_{i+1})$, $i=1,\ldots,\ell$, with $r_{\ell+1}=T$. Given the solution on $[r_{i-1},r_{i})$, with $r_{0}=0$, the value $Y_{r_{i}}$ is determined by the jump of $\mathbf{X}$ at time $r_{i}$. We therefore consider $(y_{i},y_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}];\mathbb{R}^{k})$, where
@@
for $i=\ell$. Analogously to Lemma 3.4, we can show that the functional satisfies Assumption 2.4 (i) and (ii) with constants depending additionally on $\ell$, that is $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends on $p$ and $\ell$, see the definition of $(\alpha_{i},\alpha_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}])$, and $C_{F,K,X,X}$ depends on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, and $\|X\|_{p}$. Note that it is again sufficient to check Assumption 2.4 (ii) for $X=\widetilde{X}$. We can thus again apply Theorem 2.3 to show that there exists a unique solution to the RDE (3.3) on the interval $[r_{i},r_{i+1})$, that is
@@
For $(F,F^{\prime})$, this is shown in the proof of (i), and as we aim to obtain continuity of the solution map as a function of the initial condition $(y,y^{\prime})$, not the vector field $f$, on the interval $[0,r_{1}]$ we may consider $(F,F^{\prime})=(\widetilde{F},\widetilde{F}^{\prime})$, so, $(F-\widetilde{F},F^{\prime}-\widetilde{F}^{\prime})=0$. Theorem 2.5 now gives that
$\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,r_{1}]}+\|R^{Y}-R^{\widetilde{Y}}\|_{p,[0,r_{1}]}\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p,[0,r_{1}]},$
@@
Local estimate on $[r_{i},r_{i+1}]$, $i=1,\ldots,\ell$. To apply the continuity result presented in Theorem 2.5, we need to ensure that the functionals satisfy the required assumptions. For $(F,F^{\prime}),(\widetilde{F},\widetilde{F}^{\prime})$, this is shown in the proof of (i), and for $(F-\widetilde{F},F^{\prime}-\widetilde{F}^{\prime})$, in the proof of Corollary 3.3 (ii), where the constant $C_{F-\widetilde{F}}$ depends additionally on $\ell$. Theorem 2.5 then implies that
$\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[r_{i},r_{i+1}]}+\|R^{Y}-R^{\widetilde{Y}}\|_{p,[r_{i},r_{i+1}]}\lesssim|Y_{r_{i}}-\widetilde{Y}_{r_{i}}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p,[r_{i},r_{i+1}]},$
@@
$\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2}}\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p},$